The area of the trapezoid is 44.25 square feet find x
if the top base is 6.7 inches and the height is 5.9 what is x

Answer by checkley71(8403)

5.9*(6.7+x)/2=44.25
5.9*(6.7+x)=44.25*2
39.53+5.9x=88.5
5.9x=88.5-39.53
5.9x=48.97
x=48.97/5.9
x=8.3 is the length of the other base.
proof:
5.9(6.7+8.3)/2=44.25
5.9(15/2)=44.25
5.9*7.5=44.25
44.25=44.25